Please help, can't seem to find a solution:
The intensity of light (in foot-candle) varies inversely as the square of the distance (in feet) from the light source. The intensity of light 3 feet from the source is 72 foot-candles. How far away is the source if the intensity of light is 2 foot-candles
Thanks!

Answer by Alan3354(69443)

The intensity of light (in foot-candle) varies inversely as the square of the distance (in feet) from the light source. The intensity of light 3 feet from the source is 72 foot-candles. How far away is the source if the intensity of light is 2 foot-candles
----------
What have you tried?
=========================
The intensity of light (in foot-candle) varies inversely as the square of the distance (in feet) from the light source. The intensity of light 3 feet from the source is 72 foot-candles. How far away is the source if the intensity of light is 2 foot-candles
----
I = k/d^2 where k is a constant.
72 = k/3^2
k = 72*9 = 648
============
How far away is the source if the intensity of light is 2 foot-candles
2 = 648/d^2
d^2 = 648/2 = 324
d = 18 ft